# Automatic Graph Topology-Aware Transformer

### v1

@@ I Introduction
To reduce the cost of developing GNNs, graph neural architecture search (GNAS) techniques are proposed for automatically designing excellent model architectures [25], which provides promising directions to address the above difficulties. A typical GNAS involves exploring a unified search space describing possible architectures, evaluating their performance on downstream tasks, and using optimization algorithms to search for the best architecture [26]. GNAS faces two key technical challenges: designing the search space and search strategy. Existing search spaces can be mainly divided into the macro search space and micro search space [27]. The former mainly focuses on the design of the GNN topology. The latter focuses on discovering components in GNNs that affect graph information exchange, such as aggregation functions. Recently, Zhang et al. [28] proposed a GNAS method to find optimal graph encoding strategies and model scale of graph Transformers. However, systematic exploration of topologies and graph-aware strategies in graph Transformers is lacking.
@@
The rest of this paper is organized as follows. Section II presents related works on the graph neural architecture search and graph Transformer. In Section III, we describe the proposed EGTAS framework in detail. To validate the effectiveness and the efficiency of the EGTAS, a series of experiments are conducted in Section IV along with the ablation study. Finally, we draw conclusions and introduce potential directions for future research in Section V.
@@ II Related Work
A series of studies on GNAS have been proposed to automatically search model architectures for graph tasks in different computing scenarios. This section first reviews related work on GNAS from the perspective of search space and search strategy. Then, we give a brief overview of graph-aware strategies in graph Transformers, which are closely aligned with the technical approach presented in this paper.
@@ II-A2 Search strategy
Differentiable methods [47] construct a super-network containing all candidate operations, where each operation is considered as a probability distribution over all candidate operations. These methods use a gradient-based algorithm to iteratively optimize both architecture and weights. For instance, Zhao et al. [48] first proposed a differentiable GNAS method to search the macro-level and micro-level architecture of GNN. Qin et al. [49] developed a differentiable graph architecture search with structure optimization, enabling the joint optimization of graph architecture and structure. Wei et al. [50] presented a differentiable pooling architecture search to customize the GNN architecture for GC tasks. Wei et al. [11] adopted a differentiable approach to design the topology of GNNs, which shows state-of-the-art performance on graph learning tasks. In [51], GAUSS is proposed to solve large-scale GNAS problems by introducing a lightweight supernet and joint architecture-graph sampling.
II-B Graph-aware Strategies in Graph Transformer
@@
where Eq. (1) and Eq. (2) are the MHA component, while Eq. (3) is the FFN component. $\operatorname{SoftMax}$ represents a softmax function. $\mathbf{A}^{l}_{h}\in\mathbb{R}^{n\times n}$ and $\mathbf{O}^{l}\in\mathbb{R}^{n\times d}$ are the message-passing matrix and output of $H$ multi-head attention mechanisms, respectively. $\mathbf{Q}^{l}_{h},\mathbf{K}^{l}_{h},\mathbf{V}^{l}_{h}\in\mathbb{R}^{d\times d_{h}},\mathbf{W_{O}}^{l}\in\mathbb{R}^{(H\times d_{h})\times d},\mathbf{W}_{1}^{l}\in\mathbb{R}^{d\times d_{1}},\mathbf{b}_{1}^{l}\in\mathbb{R}^{d_{1}},\mathbf{W}_{2}^{l}\in\mathbb{R}^{d_{1}\times d},\mathbf{b}_{2}^{l}\in\mathbb{R}^{d}$ are the learnable parameters. For graph learning tasks, all node features $\mathbf{X}^{0}=\left[\mathbf{x}_{1},...,\mathbf{x}_{n}\right]$ are employed as initial input. After going through $L$ blocks, we can obtain the final high-level node representation $\mathbf{X}^{L}$. Recently, designing Transformers for graph-structured data has received considerable attention [24]. Next, we review the graph Transformer architecture from the perspective of graph-aware strategies. Typical graph-aware strategies for incorporating graph information into Transformers can be divided into three groups: 1) a combination of GNNs and Transformers (CGT), 2) positional embedding with graph information (PE), and 3) attention matrix with graph information (AM). It’s worth noting that these strategies can be flexibly combined [53, 54].
II-B1 CGT
@@
where $P_{G}\in\mathbb{R}^{n\times d_{p}}$ is the graph embedding vectors, and $f_{PE}:\mathbb{R}^{n\times d_{p}}\rightarrow{\mathbb{R}^{n\times d}}$ represents a simple feed-forward neural network to align the dimensions of vectors. Existing works adopt various methods as graph embedding vectors [59], such as Laplacian eigenvectors (LE) [60] and positional encodings based on singular value decomposition (SVD) [61]. Different from the compressed adjacency matrix in the above methods, Ying et al. [53] proposed a heuristic method (called DC in this paper) to extract structural information from the adjacency matrix, where the out-degree and in-degree of each node are used to design embedding vectors. By utilizing degree centrality, Transformers can capture key information of nodes.
II-B3 AM
@@
where $P$ is the graph information. Ying et al. [53] presented a novel spatial encoding-based attention mechanism (SE) that measures the spatial relation between nodes. To capture fine-grained structural information, Zhao et al. [62] proposed a proximity-enhanced multi-head attention mechanism (PEM). Min et al. [23] developed a masking mechanism-based graph encoding strategy (GMT), where the graph-masking mechanism enforces heads to attend to different subspaces, thereby improving the representative capability of models.
@@ II-C Discussion
III EGTAS: Evolutionary Graph Transformer Architecture Search
@@
III-A Problem Definition
Definition III.1
@@ III-A Problem Definition
III-B Comprehensive Graph Transformer Search Space
@@
III-B1 Macro Search Space
@@
III-B2 Micro Search Space
@@
III-C Search Strategy
Given the comprehensive search space in Section III.B, a large number of operation combinations need to be explored by a search strategy. As indicated by Definition III.1, direct acquisition of gradient information pertaining to upper-level objectives (such as #Acc) concerning the architecture is infeasible. Therefore, GNAS is usually regarded as a black-box optimization problem [25]. Due to excellent parallelism and gradient-free properties, evolutionary search is widely employed to solve black-box optimization problems [41]. Next, we present the proposed surrogate-assisted evolutionary search from encoding strategy, surrogate model construction, and evolutionary search.
@@ III-C1 Encoding Strategy
III-C2 Surrogate Model Construction
@@ IV Experiments
IV-A Node Classification Task
IV-A1 Experimental Settings
@@ IV-A2 Experimental Results
IV-B Graph Classification Task
@@
Three state-of-the-art baselines are considered for the graph classification task: GIN [67], Graphormer [53], and AutoGT [28]. GIN and Graphormer are the popular human-designed GNN and Transformer models, respectively. AutoGT is a state-of-the-art GNAS approach to search for graph Transformer architectures for graph classification tasks, which follows the basic framework of evolutionary search and supernet training (ST). The algorithm configuration of the proposed EGTAS is consistent with Section IV.A. We employ ten-fold cross-validation for all baselines. For a fair comparison, all hyper-parameter and training strategy configurations are kept the same as the original paper of AutoGT [28].
@@ IV-C Benchmarking GNNs
We compare EGTAS with popular manual baselines: GCN [63], GAT [65], GIN [67], GatedGCN [79], PNA [80], DGN [81], CIN [82], CRaWl [83], GIN-AK+ [84], SAN [59], Graphormer [53], K-Subtree SAT [56], EGT [61], and GPS [54]. These baselines cover a variety of well-known message-passing neural networks and graph transformer models, which are reported in [54]. All baseline results are based on ten independent runs. The algorithm configuration for EGTAS aligns with the description in Section IV.A. The training hyperparameters for benchmark graph datasets are presented in Tables X.
@@ V Conclusion
In the future, we will focus on multiple objectives, such as the number of parameters and latency, to recommend practical graph Transformers for different computing scenarios. Due to the complexity of graph-structured data, Transformer architecture search on other graph tasks is also a promising direction, such as link prediction and recommendation [85]. In addition, graph learning with heterophily is receiving increasing attention [86]. Searching Transformer architectures specifically tailored for heterophilic graphs merits deeper investigation. As is often the case with GNAS, surrogate model construction on large-scale graphs suffers from computational bottlenecks. Evaluating multiple sampling architectures in parallel offers an efficient solution. Implementing the parallelization of EGTAS remains a topic for future work.

### v2

@@ I Introduction
To reduce the cost of developing GNNs, graph neural architecture search (GNAS) techniques are proposed for automatically designing excellent model architectures [25], which provides promising directions to address the above difficulties. A typical GNAS involves exploring a unified search space describing possible architectures, evaluating their performance on downstream tasks, and using optimization algorithms to search for the best architecture [26]. GNAS faces two key technical challenges: designing the search space and search strategy. Existing search spaces can be mainly divided into the macro search space and micro search space [27]. The former mainly focuses on the design of the GNN topology. The latter focuses on discovering components in GNNs that affect graph information exchange, such as aggregation functions. Recently, Zhang et al. [28] proposed AutoGT to find optimal graph encoding strategies and model scale of graph Transformers. However, systematic exploration of topologies and graph-aware strategies in graph Transformers is lacking.
@@
The rest of this paper is organized as follows. Section II presents related works on the graph neural architecture search and graph Transformer. The working mechanism of the graph Transformer is introduced in Section III. In Section IV, we describe the proposed EGTAS framework in detail. To validate the effectiveness and the efficiency of the EGTAS, a series of experiments are conducted in Section V along with the ablation study. Finally, we draw conclusions and introduce potential directions for future research in Section VI.
@@ II Related Work
A series of studies on GNAS have been proposed to automatically search model architectures for graph tasks in different computing scenarios. We first review related work on GNAS from the perspective of search space and search strategy. Then, we introduce the key techniques in graph Transformer.
@@ II-A2 Search strategy
Differentiable methods [47] construct a super-network containing all candidate operations, where each operation is considered as a probability distribution over all candidate operations. These methods use a gradient-based algorithm to iteratively optimize both architecture and weights. For instance, Zhao et al. [48] first proposed a differentiable GNAS method, SANE, to search the macro-level and micro-level architecture of GNN. Qin et al. [49] developed a differentiable graph architecture search with structure optimization, enabling the joint optimization of graph architecture and structure. Wei et al. [50] presented a differentiable pooling architecture search to customize the GNN architecture for GC tasks. Wei et al. [11] proposed F2GNN to design the topology of GNNs, which shows state-of-the-art performance on graph learning tasks. In [51], GAUSS is proposed to solve large-scale GNAS problems by introducing a lightweight supernet and joint architecture-graph sampling.
II-B Graph Transformer
Designing Transformers for graph-structured data has received considerable attention [24], aiming to alleviate the over-smoothing issue [52]. Several studies have introduced graph information into positional encodings and attention matrices, allowing Transformers to understand graph structures. SAN [53] designs a learnable positional encoding by utilizing the full Laplacian spectrum and incorporates real and virtual edge information into the attention mechanism. EGT [54] presents a positional encoding based on the singular value decomposition (SVD) and proposes edge channels to focus on edge embeddings. Graphormer [55] encodes various graph structure information into Transformer, exhibiting excellence in large-scale graph classification. GMT [23] explores graph masking mechanisms in the attention matrix, considering different types of interaction graphs. K-Subtree SAT [56] extracts a subgraph representation rooted at each node through existing GNNs before applying the attention mechanism. Rampášek et al. [57] proposed a modular framework GPS that supports multiple types of graph encodings, showing competitive results in various benchmarks. In addition, Kong et al. [58] presented a global Transformer to address node classification tasks on large-scale graphs.
III Preliminaries
@@
where Eq. (1) and Eq. (2) are the MHA component, while Eq. (3) is the FFN component. $\operatorname{SoftMax}$ represents a softmax function. $\mathbf{A}^{l}_{h}\in\mathbb{R}^{n\times n}$ and $\mathbf{O}^{l}\in\mathbb{R}^{n\times d}$ are the message-passing matrix and output of $H$ multi-head attention mechanisms, respectively. $\mathbf{Q}^{l}_{h},\mathbf{K}^{l}_{h},\mathbf{V}^{l}_{h}\in\mathbb{R}^{d\times d_{h}},\mathbf{W_{O}}^{l}\in\mathbb{R}^{(H\times d_{h})\times d},\mathbf{W}_{1}^{l}\in\mathbb{R}^{d\times d_{1}},\mathbf{b}_{1}^{l}\in\mathbb{R}^{d_{1}},\mathbf{W}_{2}^{l}\in\mathbb{R}^{d_{1}\times d},\mathbf{b}_{2}^{l}\in\mathbb{R}^{d}$ are the learnable parameters. For graph learning tasks, all node features $\mathbf{X}^{0}=\left[\mathbf{x}_{1},...,\mathbf{x}_{n}\right]$ are employed as initial input. After going through $L$ blocks, we can obtain the final high-level node representation $\mathbf{X}^{L}$.
Next, we introduce the working mechanism of graph Transformers from the perspective of graph-aware strategies, which are closely aligned with the proposed EGTAS. Typical graph-aware strategies for incorporating graph information into Transformers can be divided into three groups: 1) a combination of GNNs and Transformers (CGT), 2) positional embedding with graph information (PE), and 3) attention matrix with graph information (AM). It’s worth noting that these strategies can be flexibly combined [55, 57].
III-A CGT
@@
where $P_{G}\in\mathbb{R}^{n\times d_{p}}$ is the graph embedding vectors, and $f_{PE}:\mathbb{R}^{n\times d_{p}}\rightarrow{\mathbb{R}^{n\times d}}$ represents a simple feed-forward neural network to align the dimensions of vectors. Existing works adopt various methods as graph embedding vectors [53], such as Laplacian eigenvectors (LE) [63] and SVD-based positional encodings [54]. Different from the compressed adjacency matrix in the above methods, Ying et al. [55] proposed a heuristic method (called DC in this paper) to extract structural information from the adjacency matrix, where the out-degree and in-degree of each node are used to design embedding vectors. By utilizing degree centrality, Transformers can capture key information of nodes.
III-C AM
@@
where $P$ is the graph information. Ying et al. [55] presented a novel spatial encoding-based attention mechanism (SE) that measures the spatial relation between nodes. To capture fine-grained structural information, Zhao et al. [64] proposed a proximity-enhanced multi-head attention mechanism (PEM). Min et al. [23] developed a masking mechanism-based graph encoding strategy, where the graph-masking mechanism enforces heads to attend to different subspaces, thereby improving the representative capability of models.
@@ III-D Discussion
IV EGTAS: Evolutionary Graph Transformer Architecture Search
@@
IV-A Problem Definition
Definition IV.1
@@ IV-A Problem Definition
IV-B Comprehensive Graph Transformer Search Space
@@
IV-B1 Macro Search Space
@@
IV-B2 Micro Search Space
@@
IV-C Search Strategy
Given the comprehensive search space in Section IV.B, a large number of operation combinations need to be explored by a search strategy. As indicated by Definition IV.1, direct acquisition of gradient information pertaining to upper-level objectives (such as #Acc) concerning the architecture is infeasible. Therefore, GNAS is usually regarded as a black-box optimization problem [25]. Due to excellent parallelism and gradient-free properties, evolutionary search is widely employed to solve black-box optimization problems [41]. Next, we present the proposed surrogate-assisted evolutionary search from encoding strategy, surrogate model construction, and evolutionary search.
@@ IV-C1 Encoding Strategy
IV-C2 Surrogate Model Construction
@@ IV-C3 Evolutionary Search
IV-C4 Complexity Analysis
EGTAS consists of surrogate model construction and evolutionary search, whose time complexity is mainly determined by the training of sampling architectures (see Step 2 of Algorithm 1) and performance evaluation (see Step 12 of Algorithm 2), respectively. Assuming the number of training epochs is $N_{e}$, the size of a training set is $|\mathcal{D}_{tra}|$, and the maximum number of parameters in sampling architectures is $p$, the time complexity of training $N_{s}$ architectures is $O(N_{s}\times N_{e}\times|\mathcal{D}_{tra}|\times p)$. Evolutionary search requires $N_{p}\times T$ performance evaluations using the cheap surrogate model, where $N_{p}$ and $T$ are accordingly the population size and number of iterations. Since the training of sampling architectures is much more expensive than performance evaluation, the overall time complexity of EGTAS can be considered as $O(N_{s}\times N_{e}\times|\mathcal{D}_{tra}|\times p)$.
@@ V Experiments
V-A Node Classification Task
V-A1 Experimental Settings
@@ V-A2 Experimental Results
V-B Graph Classification Task
@@
Three state-of-the-art baselines are considered for the graph classification task: GIN [69], Graphormer [55], and AutoGT [28]. GIN and Graphormer are the popular human-designed GNN and Transformer models, respectively. AutoGT is a state-of-the-art GNAS approach to search for graph Transformer architectures for graph classification tasks, which follows the basic framework of evolutionary search and supernet training (ST). The algorithm configuration of the proposed EGTAS is consistent with Section V.A. We employ ten-fold cross-validation for all baselines. For a fair comparison, all hyper-parameter and training strategy configurations are kept the same as the original paper of AutoGT [28].
@@ V-C Benchmarking GNNs
We compare EGTAS with popular manual baselines: GCN [65], GAT [67], GIN [69], GatedGCN [81], PNA [82], DGN [83], CIN [84], CRaWl [85], GIN-AK+ [86], SAN [53], Graphormer [55], K-Subtree SAT [56], EGT [54], and GPS [57]. These baselines cover a variety of well-known message-passing neural networks and graph transformer models, which are reported in [57]. All baseline results are based on ten independent runs. The algorithm configuration for EGTAS aligns with the description in Section V.A. The training hyperparameters for benchmark graph datasets are presented in Tables X.
@@ VI Conclusion
In the future, we will focus on multiple objectives, such as the number of parameters and latency, to recommend practical graph Transformers for different computing scenarios. Due to the complexity of graph-structured data, Transformer architecture search on other graph tasks is also a promising direction, such as link prediction and recommendation [87]. In addition, graph learning with heterophily [88] or homophily [89] is receiving increasing attention. The exploration of Transformer architectures tailored to special graphs merits deeper investigation. As is often the case with GNAS, surrogate model construction on large-scale graphs suffers from computational bottlenecks. Evaluating multiple sampling architectures in parallel offers an efficient solution. Implementing the parallelization of EGTAS remains a topic for future work.